40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 4 19:31:17 2011, kiwamu, Update, SUS, PRM and BS oplev laser died oplevs.png
    Reply  Wed Oct 5 11:33:25 2011, steve, Update, SUS, PRM and BS oplev laser replaced 
       Reply  Mon Apr 20 16:42:48 2015, steve, Update, SUS, PRM and BS oplev laser replaced newOplevLaser.png
          Reply  Fri Apr 24 23:40:15 2015, rana, Update, SUS, PRM and BS oplev laser replaced BSOL.pdf
             Reply  Sun Apr 26 15:34:34 2015, Jenne, Update, SUS, PRM oplev centered 
             Reply  Tue Jun 9 10:57:12 2015, Steve, Update, SUS, BS oplev laser tuned 100dBStrend0425.png18dBSdrift.png
Message ID: 11246     Entry time: Fri Apr 24 23:40:15 2015     In reply to: 11227     Reply to this: 11256   11349
Author: rana 
Type: Update 
Category: SUS 
Subject: PRM and BS oplev laser replaced 

Recently, Steve replaced the HeNe which was sourcing the BS & PRM OL. After replacement, no one checked the beam sizes and we've been living with a mostly broken BS OL. The beam spot on the QPD was so tiny that we were seeing the 'beam is nearly the size of the segment gap' effect.

Today I removed 2 of the lenses which were in the beam path: one removed from the common PRM/BS path, and one removed from the PRM path. The beams on both the BS & PRM got bigger. The BS beam is bigger by a factor of 7. I've increased the loop gains by a factor of 6 and now the UGFs are ~6 Hz. The loop gains were much too high with the small beam spots that Steve had left there. I would prefer for the beams to be ~1.5-2x smaller than they are now, but its not terrible.

Many of the mounts on the table are low quality and not constructed stably. One of the PRM turning mirror mounts twisted all the way around when I tried to align it. This table needs some help this summer.

In the future: never try locking after an OL laser change. Always redo the telescope and alignment and check the servo shape before the OL job is done.

Also, I reduced the height of the RG3.3 in the OL loops from 30 to 18 dB. The BS OL loops were conditionally stable before and thats a no-no. It makes it oscillate if it saturates.

Attachment 1: BSOL.pdf  11 kB  | Hide | Hide all
BSOL.pdf
ELOG V3.1.3-